OpenStudy (anonymous):

what is an equation of the line that passes through the point (3,-8) and has slope -2?

OpenStudy (anonymous):

Slope of a line is y-y1=m(x-x1) m=rise/run=y2-y1/x2-x1=slope

OpenStudy (anonymous):

(x1,y1) is your point. You just plug in and solve for y=